jagr2713 (jagr2713):

f(x) =-4x3 + 7x Determine if its Even,odd, neither

OpenStudy (tennis5518):

What do u think?

jagr2713 (jagr2713):

don't we plug -x for x?

OpenStudy (freckles):

yep

OpenStudy (tennis5518):

yes

OpenStudy (tennis5518):

so what do u think now

jagr2713 (jagr2713):

so we get umm -4(-x)^3+7(-x) so neither?

OpenStudy (freckles):

it maybe help you to know (-x)^odd =-x^odd while (-x)^even=x^even

jagr2713 (jagr2713):

so it will be Odd?

OpenStudy (tennis5518):

Like what he says even thought i was going to say

OpenStudy (tennis5518):

yes

jagr2713 (jagr2713):

How do you know when its neither ?

OpenStudy (freckles):

when you plug in -x you get neither f(x) or -f(x) back.

jagr2713 (jagr2713):

huh?

OpenStudy (freckles):

definitions if f(-x)=f(x) then f is even if f(-x)=-f(x) then f is odd so if f(-x) gives you neither f(x) or -f(x) then f is neither odd or even

OpenStudy (freckles):

short cut methods for polynomials: if all powers are odd, then the function is odd if all powers are even, then function is even

OpenStudy (freckles):

notice here your -4x^3+7x is polynomial and the powers are 3 and 1 the powers are both odd so f is odd this isn't using definition it is a short cut way

OpenStudy (freckles):

another example -4x^3+5 is neither because the powers are 3 and 0 you have a mixture which means it is neither

OpenStudy (freckles):

another example -4x^2+5 is even because the powers are 2 and 0 both powers are even so the function is even

jagr2713 (jagr2713):

OHOHOHOHHO you explained it so well :D

OpenStudy (freckles):

that is a short cut method for polynomials though if the teacher ask you to show work you will need to do the definition

OpenStudy (freckles):

also you do not want to forget the definition because it will be useful for non-polynomials

jagr2713 (jagr2713):

-3x3 + 9x2 - 3 so this it odd?

OpenStudy (freckles):

I could give you a lot of short cuts for other functions but it is just too many to list and so that is why learning the definition is also good

OpenStudy (freckles):

no

OpenStudy (freckles):

you have a mixture so it is neither

OpenStudy (freckles):

you have the powers 3,2, and 0

jagr2713 (jagr2713):

Wait no neither

jagr2713 (jagr2713):

@freckles

OpenStudy (freckles):

yes?

jagr2713 (jagr2713):

so neither?

OpenStudy (freckles):

Ok I guess you didn't read my post above "you have a mixture so it is neither'

OpenStudy (freckles):

the mixture is pertaining to you have both odd and even powers in the poly

jagr2713 (jagr2713):

so 5x^2-5 is even

OpenStudy (freckles):

right because you have 2 and 0

OpenStudy (tennis5518):

ok honestly all you have to say if it wrong or right

jagr2713 (jagr2713):

Thanks freckles :D

OpenStudy (freckles):

are you talking to me @Tennis5518 ?

OpenStudy (freckles):

I think me copying and pasting what I said above was confirmation enough @Tennis5518

OpenStudy (freckles):

and np @jagr2713 how far do you want to go in math?

jagr2713 (jagr2713):

CALC!!!!!

OpenStudy (freckles):

just asking because we need to work on using the definition if you do go further

jagr2713 (jagr2713):

oH, yea some stuff i dont know lol

OpenStudy (freckles):

\[f(x)=\frac{x^2+1}{x^3-x}\] for example this is not a polynomial how would you determine if this was even or odd

OpenStudy (tennis5518):

And yes im talkin about u

jagr2713 (jagr2713):

umm idk @freckles

OpenStudy (freckles):

ok @Tennis5518 as I said I think copying and pasting what I said above was enough

OpenStudy (tennis5518):

im gave you a medal cause you deserve it but that to confusing

OpenStudy (tennis5518):

to the people that doesnt know

OpenStudy (freckles):

anyways.... I don't know what you are talking about... I'm going to input -x in place of x \[f(x)=\frac{x^2+1}{x^3-x} \\ f(-x)=\frac{(-x)^2+1}{(-x)^3-(-x)}=\frac{x^2+1}{-x^3+x}=\frac{x^2+1}{-(x^3-x)}=-\frac{x^2+1}{x^3-x}=-f(x)\] do you see what has happened here ?

OpenStudy (freckles):

like first I input -x I used what I said earlier (-x)^even=x^even and (-x)^odd=-x^odd

OpenStudy (freckles):

like (-x)^2=x^2 and (-x)^3=-x^3

OpenStudy (freckles):

now I see that the tops are already the same after doing this so I played with the bottom to see if it was the same as previous bottom or opposite

OpenStudy (tennis5518):

Ok whatever you say but im just saying your makeing it complicate it

OpenStudy (tennis5518):

making*

jagr2713 (jagr2713):

Oh i get it. Thanks :D

OpenStudy (freckles):

anyways I will stop if you think you are learning too much ? (Doctor Tennis orders :p )

jagr2713 (jagr2713):

Yea i kinda am haha

OpenStudy (freckles):

k